Consider the following.
Government spending = R550
Exports=R330
Autonomous consumption =R280
Autonomous imports =R170
Investment expenditure=R120
Marginal propensity to consume = 0,75
Full employment level of income =R5700

A. Autonomous expenditure is equal to (1)R900 (2)R960 (3)R1040 (4)R1110
B. What is the marginal propensity to save? (1)1 (2)1.60 (3)0.25 (4)0.40
C. What is the value of the multiplier? (1)2 (2)2.50 (3)4 (4)3
D. What is the equilibrium level of income Y*? (1)R5120 (2) R4440 (3)R4400 (4) R5260
1
Expert's answer
2020-04-09T08:43:36-0400

A. (4) R1110.

B.(3)0.25

"MPS=1-MPC"

"MPS = 1 - 0.75 = 0.25"

C.(3) 4.

"Multiplier = 1\/MPS=1\/0.25 = 4"

D. (2,3) R4440

"Y' = Multiplier*Autonomous spending = 4*1110 = 4440."